# High-Dimensional Bayesian Model Comparison in Cosmology with GPU-accelerated Nested Sampling and Neural Emulators

### v1

@@ 2.4 CMB TT Power Spectrum
The CMB temperature auto-power spectrum analysis is kept deliberately simple by assuming a cosmic variance-limited scenario. In this idealized case, the uncertainty in the power spectrum measurements $C_{\ell}^{\text{obs}}$ at each multipole $\ell$ is due only to the inherent statistical fluctuations of the CMB itself, rather than instrumental noise or foreground effects. The likelihood for the observed power spectrum, given a theoretical model $C_{\ell}(\theta)$ follows a chi-squared distribution of shape $2\ell+1$. The CosmoPower-JAX emulator suite provides the predictions for $C^{\rm{TT}}_{\ell}(\theta)$, allowing them to be vectorized efficiently. This likelihood is run on mock data generated from the Planck 2018 best-fit parameters (Planck Collaboration et al., 2020).
@@ 2.5 Cosmic Shear
We use the likelihood framework described in Piras and Spurio Mancini (2023); Piras et al. (2024). Cosmic shear refers to the weak gravitational lensing effect where the observed shapes of distant galaxies are coherently distorted by the intervening large-scale structure of the Universe (Kilbinger, 2015). This distortion is quantified by the angular shear power spectra $C_{\ell}^{\epsilon\epsilon}$.
The theoretical prediction for these $C_{\ell}^{\epsilon\epsilon}$ values involves integrating over the matter power spectrum $P_{\delta\delta}(k,z)$ and the lensing efficiency, which depends on the redshift distribution of source galaxies. CosmoPower-JAX is used to emulate the linear and non-linear matter power spectrum, and the $C_{\ell}^{\kappa\kappa}$ are calculated using the extended Limber approximation (LoVerde and Afshordi, 2008), which translates between Fourier modes $k$ and multipoles $\ell$. Full details of the likelihood implementation and nuisance parameters are given in appendix A.
@@
Table 2: The evidence and computation times for the models and samplers considered, where “wall-clock” and “total computation” times are quoted as the time to run both cosmologies. (I) denotes the interpolated likelihoods, and (*) denotes the times reported by reported by Piras et al. (2024). The (*) models were both run on a jax_cosmo implementation of the “full” shear likelihood.
Model Method $\log\mathcal{Z}_{\Lambda\text{CDM}}$ $\log\mathcal{Z}_{w_{0}w_{a}}$ $\log B_{01}$ Total Time Hardware
@@
Cosmic Shear GPU-NS $8.09\pm 0.31 40955.49±0$$5.49\pm 0.33$ $2.60\pm 0.47$ $\sim$2 Days 1 A100 GPU
Cosmic Shear (I) GPU-NS $40959.03\pm 0.31$ $40956.28\pm 0.32$ $2.74\pm 0.44$ $\sim$11 hours 1 A100 GPU
NUTS + harmonic $40956.67\pm 0.38$ $40954.05\pm 0.39$ $2.62\pm 0.55$ $\sim$6 Hours 1 A100 GPU
Cosmic Shear (*) polychord $107.03\pm 0.27$ $107.81\pm 0.74$ $0.78\pm 0.79$ $\sim$8 Months 48 CPUs
NUTS + harmonic $40956.55\pm 0.06$ $40955.03\pm 0.04$ $1.53\pm 0.07$ $\sim$2 Days 12 A100 GPUs
To fully leverage the vectorization capabilities of JAX, significant modifications were made to the original likelihood code. The original implementation using JAX-cosmo (Campagne et al., 2023), built around ‘cosmology’ class objects, offers high interpretability but is not amenable to JAX’s just-in-time (JIT) compilation of pure functions. We therefore refactored the likelihood into a series of stateless functions. This process resulted in a less general but highly optimized code path. Extensive validation was performed to ensure consistency between the two versions, with minor ($0.1\%$) numerical differences owing to slightly different choices of $\chi,D(\chi)$ implementations. By re-running HMC on our optimised likelihood we isolate the performance of each sampling method on the same likelihood function.
As shown in Fig. 1(b), the full Cosmic Shear likelihood calculation is compute-intensive and does not scale efficiently beyond a small number of parallel evaluations. To explore the effects of lightening this bottleneck, we also implement an alternative version of the likelihood, where the matter power spectrum emulator is evaluated over a smaller $z$-grid and interpolated. This trades accuracy for computational throughput and offers a comparison to the performance of a less computationally intensive likelihood. We report these results as the “Full” and “Interpolated” likelihoods. The same interpolated likelihood was used for the HMC comparison.
3 Results
We present the results of our comparative analyses in Table 2, detailing the Bayesian evidence values and computation times for both the CMB and Cosmic Shear problems. For each problem, we compare our GPU-accelerated Nested Sampler (GPU-NS) against a traditional CPU-based NS implementation (polychord) and a modern HMC sampler coupled with a learned harmonic mean estimator (harmonic). We ran the GPU-NS with 1000 live points in both cases, and for HMC we ran 60 chains with 100/400 warm-up/samples and 400/2000 for the CMB and shear likelihoods respectively. For the shear likelihood these numbers were chosen to match the sampling strategy of Piras et al. (2024), and for the CMB the total number of samples was scaled down for the dimension of the problem, in line with the decrease in nested samples taken. In both cases the HMC and GPU-NS contours are in excellent agreement.
3.1 CMB results
The CMB analysis highlights the ideal use-case for GPU-NS. As shown in Fig. 1(a), the emulator-based likelihood is highly vectorisable, and so our GPU-NS approach completes the analysis in just 12 seconds, a speed-up of nearly 300x compared to the 1 hour taken by polychord. While polychord updates a single live point per iteration, our method evolves 500 live points in parallel, fully exploiting the GPU’s architecture. This performance gain would be even more pronounced against traditional, non-emulated likelihoods based on codes like CAMB (Lewis and Challinor, 2011) or CLASS (Blas et al., 2011), which are significantly slower per evaluation.
@@ 3.2 Cosmic Shear Results
In the 37/39-dimensional Cosmic Shear analysis our GPU-NS framework transforms an analysis that is practically infeasible with CPU-based NS ($\sim$8 months reported in Piras et al., 2024) into a task that can be completed by nested sampling in approximately 2 days on a single A100 GPU. The cosmological posteriors are shown in Fig. 2, and the full 39-dimensional posterior is shown in appendix B. However, in contrast to the massive parallelism of the CMB-only analysis, Fig. 1(b) demonstrates that the more computationally intensive shear likelihood inhibits the potential speed-up achievable on a single GPU. Here HMC overtakes nested sampling, as it can use gradient information to explore the higher dimensional landscape just as well as the CMB likelihood.
@@ 4 Discussion
Figure 2: Marginalized posterior distributions for the cosmological parameters of the Cosmic Shear analysis, using the full likelihoods with GPU-NS. The contours show the 68% and 95% credible intervals for each parameter. The dashed lines show the fiducial values that the mock data was generated with.
@@ 4.1 Performance breakdown
On the CMB analysis, where as shown in Fig. 1(a) the likelihood perfectly vectorises well beyond our chosen resolution of 1000 live points, nested sampling finds its biggest speed-up. As presented in Yallup et al. (2025), the blackjax nested sampling implementation is a well vectorised sampling procedure, and thus can fully leverage parallel evaluation of this likelihood. To gain this level of speed-up with the HMC approach would require running 500 separate chains, and the scaling properties of the likelihood suggest that running this many chains would take no longer than running fewer chains. However, due to the need for a warm-up phase, as well as the auto-correlation that would be of concern when running chains that are too short, HMC cannot utilise a vectorised likelihood to the same degree, 500 chains of 20 samples does not have the same statistical power as 10 chains of 1000 samples.
When the likelihood is the bottleneck, such as in the cosmic shear analysis, GPU-NS loses its advantage, and the two algorithms sit on a fairly equal footing. Here all of the speed-up comes from the likelihood internally vectorising, so although the samplers cannot massively vectorise to speed up the inference, individual likelihood calls are greatly sped up by having their many internal emulator calls vectorised. The interpolated likelihood appears to benefit from being able to vectorise slightly more than the full likelihood, although its speed-up is also consistent with the fact it makes fewer emulator calls per evaluation.
There are a number of potential routes to lift this bottleneck further. All the calculations presented in this work are performed in double numeric precision: whilst it is non-trivial to ensure that sufficient accuracy in result is achievable with lower precision, it is noteworthy that much of the hardware development is targetting increasingly reduced numeric precision, and this is a development that physicists need to be alert to (Jacob et al., 2017). Reducing the required numeric precision to single float numerics would greatly increase the potential to parallelise the calculations in this work. More immediately, it is also feasible to recover massive speed-ups by deploying the calculations over multiple devices/GPU clusters. The original work of  Piras et al. (2024) distributed its MCMC chains over multiple devices, thus enabling faster (in wall-time) accumulation of multiple chains. However, multiple chains in this context primarily improve the estimate of variance of the result, in contrast population level approaches such as NSS can improve total runtime (for a fixed variance) more directly using multiple devices. We leave a multi-device implementation of this pipeline to future work.
@@ 4.2 Comparison of Methods
This work shows that with GPU-acceleration nested sampling is now feasible on high dimensional analyses such as the cosmic shear analysis here. However it should be noted that although we have provided enough speed-up to explore a 39-dimensional prior, the number of likelihood evaluations nested sampling needs to converge scales very harshly with dimension (Handley et al., 2015), and as such field-level nested sampling and other ultra-high dimensional problem are still out of reach, even with these accelerations. In comparison HMC, due to its usage of gradient information, is able to explore posteriors in much higher dimensions and has already been used in field-level inference. This is a significant drawback of nested sampling, with JAX-based pipelines differentiability in our models is free information that at the present isn’t being used by nested sampling.
@@
Some caution should be taken with the error bars provided by the learned harmonic mean estimator; it represents only the scatter in the estimate taken on different subsets of the data. It does not include any error from how the flow might change if trained on a different set of posterior samples, and most importantly there is no systematic error from the flow not capturing the true posterior shape. This incomplete error budget is a concern, as presented the GPU-NS versus HMC results in table 2 are around $4\sigma$ away from each-other for the interpolated likelihood, and while each method appears to recover the same Bayes factor there is no guarantee this would be true in general, and a discrepancy of 2.5 in log space is the difference between decisive and inconclusive evidence when performing Bayesian model comparison. It should be noted the discrepancies between the Bayes factors of this work and of Piras et al. (2024) may have arisen due to the minor numerical differences in the likelihood implementations mentioned before, so cannot be compared as directly as our HMC (interpolated) versus GPU-NS (interpolated) results.
@@
Nested Sampling, in turn, is not immune to its own implementation-specific biases, and its error bar is also conditional. The NS error calculation assumes that at each iteration, a new live point is drawn perfectly uniformly from the prior volume constrained by the current likelihood threshold (Skilling, 2006). In practice, this is approximated with an inner MCMC sampler which, in high-dimensional or complex posteriors, can struggle to explore the entire valid region. This inefficiency can introduce a systematic bias that the reported statistical error does not account for. However, since nested sampling is a global algorithm (sampling over the entire prior instead of rapidly settling into the posterior), it can properly account for the error in the specific path taken when crossing from the prior to the posterior, and it is unlikely to miss modes when run at sufficient precision. The secondary $A_{\text{IA},10}$ peak, for example, was found in very few of the HMC chains, due to the sampler’s sensitivity to its initialisation, and was thus under-represented by those posterior samples. We stress that this is an unphysical mode in a nuisance parameter, and may well be just an artifact of the formulation of the likelihood.
@@
Bringing the overall runtime of this challenging frontier problem to a level that multiple methods can be tested with rapid turn around is a significant success of the optimizations performed as part of this work. It is now possible to investigate the discrepancy between the harmonic and NS based pipelines by incorporating other advanced MCMC based normalizing constant estimators such as Sequential Monte Carlo (Doucet et al., 2001; Chopin and Papaspiliopoulos, 2020).
@@ 5 Conclusion
We have demonstrated that the combination of GPU-acceleration, JAX-based emulators, and a vectorized Nested Sampling algorithm removes the primary computational barrier in using Nested Sampling in high-dimensional problems. Our framework achieves a speed-up of over two orders of magnitude on a cosmic-variance-only CMB analysis and, more critically, reduces the runtime of a 39-dimensional cosmic shear analysis from months to a matter of days, placing Nested Sampling on an equal computational footing with the fastest alternative methods.
@@
Nested sampling is uniquely positioned within the field of cosmology, as a reliable baseline for accurate evidence calculation; however, its implementation is typically tied to legacy code pipelines. This work establishes that, through the use algorithms and forward models built to leverage modern hardware, nested sampling can be a routine component of the next generation of cosmological analyses. As we move towards future experiments, we must be mindful of trading off precision in order to make analyses computationally feasible, and this work demonstrates that nested sampling will continue to be a viable tool in the next stage of cosmological surveys and experiments.
While this and other work (Yallup et al., 2025; Prathaban et al., 2025) establishes GPU nested sampling as being able to greatly accelerate medium to high dimensional analysis, future work will push the scale of feasible analyses even further, now made possible by building our inference pipelines around GPU hardware. Developments such as efficient parallelisation of multi-GPU instances, as well as efforts at pushing emulation even further, will enable an acceleration of the next tier of problems.

### v2

@@ 2.4 CMB TT Power Spectrum
The CMB temperature auto-power spectrum analysis is kept deliberately simple by assuming a cosmic variance-limited scenario. In this idealised case, the uncertainty in the power spectrum measurements $C_{\ell}^{\text{obs}}$ at each multipole $\ell$ is due only to the inherent statistical fluctuations of the CMB itself, rather than instrumental noise or foreground effects. The likelihood for the observed power spectrum, given a theoretical model $C_{\ell}(\theta)$ follows a chi-squared distribution of shape $2\ell+1$. The CosmoPower-JAX emulator suite provides the predictions for $C^{\rm{TT}}_{\ell}(\theta)$, allowing them to be vectorised efficiently. This likelihood is run on mock data generated from the Planck 2018 best-fit parameters (Planck Collaboration et al.,, 2020).
@@ 2.5 Cosmic Shear
We use the likelihood framework described in Piras and Spurio Mancini, (2023); Piras et al., (2024). Cosmic shear refers to the weak gravitational lensing effect where the observed shapes of distant galaxies are coherently distorted by the intervening large-scale structure of the Universe (Kilbinger,, 2015). This distortion is quantified by the angular shear power spectra $C_{\ell}^{\epsilon\epsilon}$.
The theoretical prediction for these $C_{\ell}^{\epsilon\epsilon}$ values involves integrating over the matter power spectrum $P_{\delta\delta}(k,z)$ and the lensing efficiency, which depends on the redshift distribution of source galaxies. CosmoPower-JAX is used to emulate the linear and non-linear matter power spectrum, and the $C_{\ell}^{\epsilon\epsilon}$ are calculated using the extended Limber approximation (LoVerde and Afshordi,, 2008), which translates between Fourier modes $k$ and multipoles $\ell$. Full details of the likelihood implementation and nuisance parameters are given in Appendix A.
@@
To fully leverage the vectorisation capabilities of JAX, significant modifications were made to the original likelihood code. The original implementation using JAX-cosmo (Campagne et al.,, 2023), built around ‘cosmology’ class objects, offers high interpretability but is not amenable to JAX’s just-in-time (JIT) compilation of pure functions. We therefore refactored the likelihood into a series of stateless functions. This process resulted in a less general but highly optimised code path. Extensive validation was performed to ensure consistency between the two versions, with minor ($0.1\%$) numerical differences owing to slightly different choices of $\chi$ and $D(\chi)$ implementations. By re-running HMC on our optimised likelihood we isolate the performance of each sampling method on the same likelihood function.
As shown in Fig. 1(b), the full cosmic shear likelihood calculation is compute-intensive and does not scale efficiently beyond a small number of parallel evaluations. To explore the effects of alleviating this bottleneck, we also implement an alternative version of the likelihood, where the matter power spectrum emulator is evaluated over a smaller $z$-grid and interpolated. This trades accuracy for computational throughput and offers a comparison to the performance of a less computationally intensive likelihood. We report these results as the “Full” and “Interpolated” likelihoods.
3 Results
We present the results of our comparative analyses in Table 2, detailing the Bayesian evidence values and computation times for both the CMB and cosmic shear problems. For each problem, we compare our GPU-accelerated Nested Sampler (GPU-NS) against a traditional CPU-based NS implementation (polychord) and a modern HMC sampler coupled with a learned harmonic mean estimator (harmonic). We ran the GPU-NS with 1000 live points in both cases, and for HMC we ran 60 chains with 100/400 warm-up/samples and 400/2000 for the CMB and shear likelihoods respectively. This choice of 60 chains of 2000 samples, as well as all other tuning details of the HMC implementation, replicates Piras et al., (2024). In both cases the HMC and GPU-NS contours are in excellent agreement.
3.1 CMB results
Table 2: The evidence and computation times for the models and samplers considered, where “wall-clock” times and “GPU-hours” are quoted as the time to run both cosmologies, and are given by their scale (days, hours, minutes) for interpretability. (I) denotes the interpolated likelihoods, and (*) denotes the times reported by Piras et al., (2024). The (*) models were both run on a jax_cosmo implementation of the “full” shear likelihood, where the polychord likelihood differs by a normalisation factor, as well as using CAMB instead of emulation for calculating the matter power spectrum. In this table, the GPU-hours and $\log B_{01}$ columns provide the clearest comparison between the NUTS+harmonic and GPU-NS methods.
Model Method $\log\mathcal{Z}_{\Lambda\text{CDM}}$ $\log\mathcal{Z}_{w_{0}w_{a}}$ $\bm{\log B_{01}}$ Wall Clock Hardware GPU-Hours
@@ 3.1 CMB results
Shear GPU-NS $8.09\pm 0.31 40955.49±0$$5.49\pm 0.33$ $2.60\pm 0.47$ 2 Days 1 A100 GPU 48
NUTS + harmonic $40956.58\pm 0.14$ $40954.38\pm 0.31$ $2.20\pm 0.34$ 10 Hours 1 A100 GPU 10
Shear (I) GPU-NS $40959.03\pm 0.31$ $40956.28\pm 0.32$ $2.74\pm 0.44$ 11 hours 1 A100 GPU 11
NUTS + harmonic $40956.67\pm 0.38$ $40954.05\pm 0.39$ $2.62\pm 0.55$ 6 Hours 1 A100 GPU 6
Shear (*) polychord + CAMB $107.03\pm 0.27$ $107.81\pm 0.74$ $0.78\pm 0.79$ 8 Months 48 CPUs -
NUTS + harmonic $40956.55\pm 0.06$ $40955.03\pm 0.04$ $1.53\pm 0.07$ 2 Days 12 A100 GPUs 576
The CMB analysis highlights the ideal use-case for GPU-NS. As shown in Fig. 1(a), the emulator-based likelihood is highly vectorisable, and so our GPU-NS approach completes the analysis in just 12 seconds, a speed-up of nearly 300x compared to the 1 hour taken by polychord. While polychord updates a single live point per iteration, our method evolves 500 live points in parallel, fully exploiting the GPU’s architecture. This performance gain would be even more pronounced against traditional, non-emulated likelihoods based on codes like CAMB (Lewis et al.,, 2000; Lewis and Challinor,, 2011; Howlett et al.,, 2012) or CLASS (Lesgourgues,, 2011; Lesgourgues and Tram,, 2011; Blas et al.,, 2011), which are significantly slower per evaluation.
@@ 3.2 Cosmic Shear Results
In the 37/39-dimensional cosmic shear analysis our GPU-NS framework transforms an analysis that is practically infeasible with CPU-based NS ($\sim$8 months reported in Piras et al.,, 2024) into a task that can be completed by nested sampling in approximately 2 days on a single A100 GPU. The cosmological posteriors are shown in Fig. 2, and the full 39-dimensional posterior is shown in Appendix C. However, in contrast to the massive parallelism of the CMB-only analysis, Fig. 1(b) demonstrates that the more computationally intensive shear likelihood inhibits the potential speed-up achievable on a single GPU. Here HMC overtakes nested sampling, as it can use gradient information to explore the higher dimensional landscape just as well as the CMB likelihood.
[…]
@@ 4 Discussion
Figure 2: Marginalised posterior distributions for the cosmological parameters of the cosmic shear analysis, using the full likelihoods with GPU-NS. The contours show the 68% and 95% credible intervals for each parameter. The dashed lines show the fiducial values that the mock data was generated with.
@@ 4.1 Performance breakdown
On the CMB analysis, where as shown in Fig. 1(a) the likelihood perfectly vectorises well beyond our chosen resolution of 1000 live points, nested sampling finds its biggest speed-up. As presented in Yallup et al., (2025), the blackjax nested sampling implementation is a well-vectorised sampling procedure, and thus can fully leverage parallel evaluation of this likelihood.
While it is feasible to gain this level of speed-up with HMC on a vectorised likelihood, its statistical validity relies on using a few long chains rather than many short ones. MCMC requires chains to be sufficiently long to both discard an initial warm-up sequence and to generate samples with low auto-correlation. A sampling strategy of 500 chains of 20 samples would be statistically poor; it would be dominated by warm-up overhead and the resulting chains would be too short to effectively explore the parameter space. This is in contrast to a run of 10 chains with 1000 samples each, which provides a much more robust estimate of the posterior, but is only able to calculate 10 likelihood evaluations in parallel.
When the likelihood is the bottleneck, such as in the cosmic shear analysis, GPU-NS loses its advantage, and the two algorithms sit on a fairly equal footing. Here all of the speed-up comes from the likelihood internally vectorising:although the samplers cannot massively vectorise to accelerate the inference, individual likelihood calls are greatly sped up by having their many internal emulator calls vectorised. The interpolated likelihood results are very instructive, because the change affects the speed of the GPU-NS and HMC analyses differently. While HMC runs twice as fast on the interpolated likelihood, GPU-NS sees a speed-up of more than four times. This is likely from it better utilising the increased vectorisation shown in Figure 1(b), which comes from reducing the number of emulator calls per likelihood. This, along with the CMB timing results, supports that GPU-NS is very efficient when run on vectorised likelihoods.
There are a number of potential routes to lift this bottleneck further. All the calculations presented in this work are performed in double numeric precision: whilst it is non-trivial to ensure that sufficient accuracy is achievable with lower precision, it is noteworthy that much of the hardware development is targeting increasingly reduced numeric precision, and this is a development that physicists need to be alert to (Jacob et al.,, 2017). Reducing the required numeric precision to single float numerics would greatly increase the potential to parallelise the calculations in this work. More immediately, it is also feasible to recover massive speed-ups by deploying the calculations over multiple devices/GPU clusters. The original work of Piras et al., (2024) distributed its MCMC chains over multiple devices, thus enabling faster (in wall-time) accumulation of multiple chains. However, multiple chains in this context primarily improve the estimate of the variance of the result; in contrast, population level approaches such as NSS can use multiple devices to directly improve run-time (at a fixed variance). We leave a multi-device implementation of this pipeline to future work.
@@ 4.2 Comparison of Methods
This work shows that with GPU-acceleration nested sampling is now feasible on high-dimensional analyses such as the cosmic shear analysis presented here. However, it should be noted that although we have provided enough speed-up to explore a 39-dimensional prior, the number of likelihood evaluations nested sampling needs to converge scales very harshly with dimension (Handley et al., 2015b, ), and as such field-level nested sampling and other ultra-high-dimensional problems are still out of reach, even with these accelerations. In comparison HMC, due to its usage of gradient information, is able to explore posteriors in much higher dimensions and has already been used in field-level inference. This is a significant drawback of nested sampling, with JAX-based pipelines all of our models are auto-differentiable, information that at the present is not being used by nested sampling.
@@
Some caution should be taken with the error bars provided by the learned harmonic mean estimator, as it represents only the scatter in the estimate taken on different subsets of the data. It does not include any error from how the flow might change if trained on a different set of posterior samples, and most importantly there is no systematic error from the flow not capturing the true posterior shape. This incomplete error budget is a concern: the GPU-NS versus HMC results presented in Table 2 are around $4\sigma$ away from each other for both the full and interpolated likelihoods; so while each method appears to recover the same Bayes factor, a discrepancy of 2 in log space is the difference between decisive and inconclusive evidence when performing Bayesian model comparison (Jeffreys,, 1961), so models certainly cannot be compared across methods (GPU-NS against HMC), even on identical likelihoods. It should be noted that the discrepancies between the Bayes factors of this work and of Piras et al., (2024) may have arisen due to the minor numerical differences in the likelihood implementations mentioned before, so should not be compared as directly as our HMC and GPU-NS results.
@@
Nested Sampling, in turn, is not immune to its own implementation-specific biases, and its error bar is also conditional. The NS error calculation assumes that at each iteration, a new live point is drawn perfectly uniformly from the prior volume constrained by the current likelihood threshold (Skilling,, 2006). In practice, this is approximated with an inner MCMC sampler which, in high-dimensional or complex posteriors, can struggle to explore the entire valid region. This inefficiency can introduce a systematic bias that the reported statistical error does not account for. However, since nested sampling is a global algorithm (sampling over the entire prior instead of rapidly settling into the posterior), it can properly account for the error in the specific path taken when crossing from the prior to the posterior, and it is unlikely to miss modes when run at sufficient precision. The secondary $A_{\text{IA},10}$ peak, for example, was found in very few of the HMC chains, due to the sampler’s sensitivity to its initialisation, and was thus under-represented by those posterior samples. We stress that this is an unphysical mode in a nuisance parameter, and may well be just an artefact of the formulation of the likelihood.
@@
Bringing the overall runtime of this challenging frontier problem to a level that multiple methods can be tested with rapid turn around is a significant success of the optimisations performed as part of this work. It is now possible to investigate the discrepancy between the harmonic and NS based pipelines by incorporating other advanced MCMC-based normalising constant estimators such as Sequential Monte Carlo (Doucet et al.,, 2001; Chopin and Papaspiliopoulos,, 2020).
@@ 5 Conclusion
We have demonstrated that the combination of GPU-acceleration, JAX-based emulators, and a vectorised Nested Sampling algorithm removes the primary computational barrier in using Nested Sampling in high-dimensional problems. Our framework achieves a speed-up of over two orders of magnitude on a cosmic-variance-only CMB analysis and, more critically, reduces the runtime of a 39-dimensional cosmic shear analysis from months to a matter of days, placing Nested Sampling on an equal computational footing with the fastest alternative methods.
@@
Nested sampling is uniquely positioned within the field of cosmology as a reliable baseline for accurate evidence calculation; however, its implementation is typically tied to legacy code pipelines. This work establishes that, through the use of algorithms and forward models built to leverage modern hardware, nested sampling can be a routine component of the next generation of cosmological analyses. As we move towards future experiments, we must be mindful of trading off precision in order to make analyses computationally feasible, and this work demonstrates that nested sampling will continue to be a viable tool in the next stage of cosmological surveys and experiments.
While this and other work (Yallup et al.,, 2025; Prathaban et al.,, 2025) establishes GPU nested sampling as being able to greatly accelerate medium- to high-dimensional analysis, future work will push the scale of feasible analyses even further, now made possible by building our inference pipelines around GPU hardware. Developments such as efficient parallelisation of multi-GPU instances, as well as efforts at pushing emulation even further, will enable the acceleration of the next tier of problems.